# Multi-Level Feature Aggregation and Recursive Alignment Network for Real-Time Semantic Segmentation

### v1

@@ V Conclusion
In this paper, we present a Multi-level Feature Aggregation and Recursive Alignment Network (MFARANet) to perform multi-scale inference in a single-pass network, a novel architecture for accurate real-time semantic segmentation. Specifically, our method consists of three core components: Multi-level Feature Aggregation Module (MFAM), Recursive Alignment Module (RAM) and Adaptive Scores Fusion Module (ASFM). MFAM is proposed to aggregate the hierarchical features of the encoder into four independent scales, which not only provides multi-level information for accurate segmentation, but also benefits the following spatial alignment operation. Combining the flow-based alignment module with the recursive upsampling architecture, our RAM can perform efficient and accurate spatial alignment between scaled score maps. Finally, we design ASFM to adaptively fuse multi-scale scores to generate the final prediction that favor objects of various scales. Experiments on three challenging datasets show that our MFARANet can provide a general and effective solution for real-time semantic segmentation.

### v3

@@ 6 Conclusion
In this paper, we present a Multi-level Feature Aggregation and Recursive Alignment Network (MFARANet) to perform multi-scale parallel inference in a single-pass network, a novel architecture for accurate real-time semantic segmentation. Specifically, our method consists of three core components: Multi-level Feature Aggregation Module (MFAM), Recursive Alignment Module (RAM) and Adaptive Scores Fusion Module (ASFM). MFAM is proposed to aggregate the hierarchical features of the encoder into four independent scales, which not only provides multi-level information for accurate segmentation, but also benefits the following spatial alignment operation. Combining the flow-based alignment module with the recursive upsampling architecture, our RAM can perform efficient and accurate spatial alignment between scaled score maps. Finally, we design ASFM to adaptively fuse multi-scale scores to generate the final prediction that favor objects of various scales. Experiments on three challenging datasets show that our MFARANet can provide a general and effective solution for real-time semantic segmentation.
Statements of Ethical Approval
We used three published datasets, so ethical approval was not required.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.